MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2008/03/06
OBSERVER: Sumi Paul
CONDITION: Fine with a bit of wind
COMMENT: D3490 (60s OK), D3491 (300s OK), The telescope had a brief GPS transmission error thet halted @moa_cam3
Compressor gas pressure: 1.9 MPa
Expected Observation Time: START 21:26 END 06:34


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB11620-11641221277620
LMC2543-254973610800
SMC612-61212600
ALERT-000
GRB-000
TOTAL---3016519020 (05h17m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
55552244421444
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
44444444440


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
111111777212
Field13Field14Field15Field16
2200


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
110000000000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
21:30observingfine, light wind
22:30observingfine, light wind
23:30observingfine, light wind
00:30observingfine, light wind
01:30observingfine, light wind
02:30observingfine, light wind
03:30observingfine, light wind
04:30observingfine, light wind
05:30observingfine, light wind
06:30observingdawn is here


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day